Sol Summary- May 02nd

Crew 211 Sol 11 Summary Report 2-May-2019

Sol: 11

Summary Title: Sunset on Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The commander and vice-commander began the day with a nice morning EVA along the Hab Ridge Road. After the breakfast, the commander and vice-commander deployed their respective experiment during an engineering EVA while the rest of the team finalized their own experiments. The deployment was postponed to this morning as different minor bugs prevented us to do it yesterday evening.The GreenHab officer harvested the Spirulina that was growing since more than a week and started to dry it. Part of the crew left at 5 pm for an EVA to Grey Moon. Tonight we will all observe the stars, enjoying the clear sky that the Utah offers.

Look Ahead Plan: We planned two EVAs for tomorrow as it is our last day on Mars.

Part of the crew will go to North Ridge in the morning. The rest of the team will end their experiment and start writing the summary report. We will get back the antennas and the cubelanders, in the morning during an engineering EVA. In the afternoon, we will do a second EVA, again to North Ridge. The rest of the afternoon will be devoted to packing our cases.

Anomalies in work: No issues.

Weather: Sunny, almost no wind

Crew Physical Status: Good

EVA: The morning EVA along Hab Ridge Road was shortened by the presence of steam in our helmets due to the low temperature at this time of the day. We didn’t manage to find the sinkhole but we had nevertheless the occasion to take nice pictures with the morning light. The second EVA to Grey moon started at 5 pm. The EVA team saw breathtaking panoramas.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, EVA Report, Daily Photo Report

Support Requested: None